. Let the angle be the angle that the vector makes with
the measured counterclockwise from that axis. Find
the angle for a vector that has the following components:
(a)Ax=2.00m, Ay=-1.00m (b)Ax=2.00m, Ay=1.00m
(c)Ax= -2.00m, Ay=1.00m
1
Expert's answer
2019-10-11T10:47:16-0400

a)


"\theta =360\degree-\arctan{\frac{1}{2}}=333.4\degree"

b)


"\theta =\arctan{\frac{1}{2}}=26.6\degree"

c)


"\theta =180\degree-\arctan{\frac{1}{2}}=153.4\degree"